Chris and Pamela leave home at the same time. Chris drives eastward at a rate of 40 miles per hour, while Pamela drives south at 60 miles per hour. Find how far apart they are to the nearest mile after 3 hours.
Answer by mananth(16946)

At any given time if you draw a line bewteen chris & pamela with the starting point it will be a right triangle
d=rt
after 3 hours chris will be 120 kms from starting point
after 4 hours pamela will be 180 kms from starting point

Pythagoras theorem
The distance between them is the hypotenuse
Hypotenuse ^2= leg1^2+leg2^2

leg1 = 120 m
Leg2^2= 180 m

hypotenuse ^2= 120 ^2 + 180 ^2
hypotenuse ^2= 14400 + 32400
hypotenuse ^2= 46800
square root = 216.33 km
the distance between them after 3 hours